OpenStudy (anonymous):

a sealed flask contains 1.3 x 10^25 molecules of CO2. How many moles of CO2 are sealed in the flask?

OpenStudy (espex):

\[1.3 \times 10^{25} molecules\space CO_2 \times \frac{1 mol\space CO_2}{6.0221415 × 10^{23}\space molecules \space CO_2}=21.59mol\space CO_2\]